A device known as AtwoodÂ’s machine consists of tw
Answer #1

equations of motion for the mA is

mA*g - T = mA*a

T = mA*(g-a)............(1)


equations of motion for mB is

T-(mB*g) = mB*a

T = mB*(a+g)...(2)

(1) = (2)

mA*g - mA*a = mB*a +mB*g

accelaration a = (mA-mB)*g / (mA+mB)


and tension T = mB*g*(mA-mB+mA+mB)/(mA+mB)

T = 2*mA*mB*g/(mA+mB)